Trace the curve x3 + y2 = 2axy 


1
Expert's answer
2020-12-29T17:44:15-0500

Equation of the curve :- "\pmb {x^3+y^2=2axy}"

"x^3+y^2=2axy\\
\Rightarrow y^2-2axy+x^3=0\\
\Rightarrow y=\frac{2ax \pm \sqrt{4a^2x^2-4x^3}}{2}=ax\pm x\sqrt{a^2-x}"


"\therefore" The curve is combination of two Branches :

"y=ax+ x\sqrt{a^2-x}" ...[Branch 1]

and, "y=ax- x\sqrt{a^2-x}" ...[ Branch 2]


Differentiating w.r.t x,

"\frac{dy}{dx}=a\pm \left(\sqrt{a^2-x} -\frac{x}{2\sqrt{a^2-x} }\right)"


Symmetry :-

The curve is not symmetric w.r.t. any line or axis.


Region :-

"y=ax \pm x\sqrt{a^2-x} \\
y \in \R \Rightarrow a^2-x >0 \Rightarrow x<a^2"


"\therefore x \in (-\infty,a^2), y\in (-\infty, \infty)"


Point of Intersection :-

Putting "x=0," "y=0"

Putting "y=0," "x=0"

"\therefore" The curve passes through origin.


Let "(h,k)" be the intersection of Branch 1 and Branch 2.


"\therefore ah+h\sqrt{a^2-h}=ah-h\sqrt{a^2-h}\\
\Rightarrow h\sqrt{a^2-h}=0\\
\Rightarrow h= 0 \; or, \; a^2\\
\Rightarrow k=0 \; or, \; a^3"


So, Branch 1 and Branch 2 meet at "(0,0)" and "(a^2,a^3)".


Tangents at Origin :-

At "(0,0) , \frac{dy}{dx}=a\pm \left(\sqrt{a^2-0}-0\right)=a\pm a =2a\; or, \; 0"

"\therefore" Tangents at origin are :-

"y=2ax \; \;\;and\;\;\; y=0"


So, "(0,0) \; is \; a \; node." [ As two tangents are distinct ]


Asymptotes :-

Coefficient of highest degree of "x" and "y" are both constant. So, There does not exist any horizontal or vertical asymptote.


Putting "y=mx+c,\\"

"x^3+m^2x^2+2mcx+c^2=2axy"

"\Rightarrow" Coff. of "x^3=1=Constant"

So There does not exist any oblique asymptote either.


Derivatives :-

"y=ax\pm x\sqrt{a^2-x} \\
\Rightarrow \frac{dy}{dx}=a\pm \left(\sqrt{a^2-x} -\frac{x}{2\sqrt{a^2-x} }\right)\\"


Now, "\frac{dy}{dx}=0"

"\Rightarrow a= \pm \frac{2(a^2-x)-x}{2\sqrt{a^2-x} }\\
\Rightarrow4a^2(a^2-x)=4a^4-12a^2x+9x^2\\
\Rightarrow x=0 \; or, \; x=\frac{8a^2}{9}"


"\therefore \frac{dy}{dx}=0" at "x=8a^2\/9" for Branch 1

at "x=0" for Branch 2


Now,

for "\pmb{a>0, }"

"Branch \;1"

"y=ax+ x\sqrt{a^2-x} \\\Rightarrow \frac{d^2y}{dx^2}= \frac{3x-4a^2}{4(a^2-x)^{3/2}}"

At "x=\frac{8a^2}{9}, \; \frac{d^2y}{dx^2}=-\frac{9}{a}<0"


So, Branch 1 has a local maxima at "x=\frac{8a^2}{9}" .



"Branch \;2"


"y=ax- x\sqrt{a^2-x} \\\Rightarrow \frac{d^2y}{dx^2}= -\frac{3x-4a^2}{4(a^2-x)^{3/2}}"

At "x=0, \; \frac{d^2y}{dx^2}=\frac{1}{a}>0"


So, Branch 2 has a local minima at "x=0" .


"\pmb{(a>0)}"


for "\pmb{a<0,}"

"Branch \;1"

"y=ax+ x\sqrt{a^2-x} \\\Rightarrow \frac{d^2y}{dx^2}= \frac{3x-4a^2}{4(a^2-x)^{3/2}}"

At "x=\frac{8a^2}{9}, \; \frac{d^2y}{dx^2}=-\frac{9}{a}>0"


So, Branch 1 has a local minima at "x=\frac{8a^2}{9}" .



"Branch \;2"

"y=ax- x\sqrt{a^2-x} \\\Rightarrow \frac{d^2y}{dx^2}= -\frac{3x-4a^2}{4(a^2-x)^{3/2}}"

At "x=0, \; \frac{d^2y}{dx^2}=\frac{1}{a}<0"


So, Branch 2 has a local maxima at "x=0" .



"\pmb{(a<0)}"